OpenStudy (anonymous):

What is the 8th term of the geometric sequence 3, 9, 27, …?

OpenStudy (anonymous):

81

OpenStudy (anonymous):

\[3^8=6561\]

OpenStudy (anonymous):

how did u get that LGP?

OpenStudy (anonymous):

OHHH wait i thought it was the next sequence, srrry....u are rite

OpenStudy (anonymous):

The Geometric Progression is given to by the formula: \[a _{n}=a _{n1}r^{n-1}\] In this case, the first therm is 3, same as the r \[\frac{a_{n+1}}{a_{n}}\] So, \[a_{8}=3.3^{8-1}=3.3^7=3^8\]